OpenStudy (anonymous):

The population of honey bees is declining at a rate of 2.25% per year. If the population is 12.6 million in the Southwest, then what will the population be in two years? A. 11.5 million B. 11.78 million C. 12.04 million D. 12.13 million

OpenStudy (kropot72):

At the end of the first year the population will be \[12.6\times \frac{100-2.25}{100}=12.6\times 0.9775\ million\] At the end of the second year the population will be \[12.6\times (0.9775)^{2}\ million\]